# Adaptive Coded Federated Learning: Privacy Preservation and Straggler Mitigation

### v1

@@ V Simulations
In this section, we evaluate the performance of ACFL through simulations on a linear regression problem, as detailed in Section II. We consider an FL setting with $N=100$ devices, each possessing a local dataset with $m_{i}=100$ samples for all $i$, with $d=10$ and $o=10$. The elements of each local dataset ${{\mathbf{X}}^{(i)}}$ are i.i.d., drawn from a uniform distribution over the interval $[-1,1]$. A true parameter matrix ${{\mathbf{W}}^{True}}\in{\mathbb{R}^{d\times o}}$ is generated with elements uniformly distributed within $[0,1/30]$. The labels ${{\mathbf{Y}}^{(i)}}$ for each device are subsequently generated according to ${{\mathbf{Y}}^{(i)}}={{\mathbf{X}}^{(i)}}{{\mathbf{W}}^{True}}$, for all $i$. The initial parameter matrix ${\mathbf{W}}_{0}$ is randomly generated with elements drawn from a uniform distribution in the interval $[0,1/30]$, independent of the true parameter matrix generation.
To evaluate the effectiveness of the adaptive policy in ACFL, we conduct a comparative analysis between ACFL and its non-adaptive counterpart (denoted as NA), which employs a fixed aggregation weight $\alpha_{t}=0.5$, as suggested by [24]. In Fig. 3, we depict the training loss as a function of number of iterations for both ACFL and NA under different variances of additive noise, where the learning rate is set as $\eta_{t}=0.0001/t$ over the iterations and we consider two distinct scenarios characterized by varying straggler probabilities. As anticipated, an increase in noise variances, aimed at enhancing privacy, inversely affects the learning performance of both methods. This empirical observation aligns with our theoretical analysis, underscoring the intrinsic trade off between privacy preservation and learning performance. Notably, the performance deterioration in ACFL remains relatively modest in comparison to NA upon increasing the noise, while there is a pronounced decline in the learning performance of NA. This discrepancy underscores the superior capability of ACFL to leverage gradients from both the central server and devices more effectively with the adaptive policy and validates the advantage of employing the adaptive policy in ACFL for mitigating the adverse effects of privacy-enhancing measures on learning performance.
@@
To further validate the efficacy of the proposed method, we compare its performance with the state-of-the-art method SCFL, which generates local coded datasets by adding noise and implementing random projections. Note that the difference between SCFL and NA lies in the way of encoding local datasets, and both of them use non-adaptive aggregation weights at the central server. In Fig. 4, we plot the training loss as a function of number of iterations for both methods under various privacy levels, i.e., under varying values of $\epsilon$ using the metric of $\epsilon$-MI-DP. The variances of the noise in both methods are calculated according to Theorem 1 from this paper and Theorem 2 in [24], ensuring required value of $\epsilon$ for both methods. Here, we fix $p=0.2$ and set the learning rate as $\eta_{t}=0.0001/t$ throughout the iterations. For an equitable comparison, we conduct full-batch gradients for both methods [^1] It is important to note that the analyses and simulations presented in this paper can be readily extended to scenarios involving stochastic gradient computations by the devices and the central server.. Furthermore, the local coded datasets in both methods are of the same size. From Fig. 4, it is evident that ACFL achieves superior learning performance under equivalent privacy level. Moreover, the learning performance of ACFL exhibits greater resilience and suffers less from increasing the noise variances to heighten the privacy level. This advantage stems from the adoption of an adaptive policy in ACFL for determining optimal aggregation weights, thereby facilitating better performance of privacy and learning.
@@
Figure 4: Training loss as a function of number of iterations of ACFL and SCFL under varying privacy variances.
@@ VI Conclusions
In this article, we addressed the problem of FL in the presence of stragglers. For this issue, the CFL paradigm was explored in the literature. However, the aggregation of gradients at the central server in existing CFL methods, which employ fixed weights across iterations, leads to diminished learning performance. To overcome this limitation, we proposed a new method: ACFL. In ACFL, each device uploads a coded local dataset with additive noise to the central server before training, enabling the central server to generate a global coded dataset under privacy preservation requirements. During each iteration of the training, the central server aggregates the received gradients as well as the gradient computed from the global coded dataset by using an adaptive policy to determine the aggregation weights. This approach optimizes the performance in terms of privacy and learning through convergence analysis and MI-DP. Finally, we provided simulation results to demonstrate the superiority of ACFL compared to non-adaptive methods. It is worth noting that in this work, we have focused on the privacy preservation of raw datasets of edge devices. In the future, we plan to extend the proposed method to more complex scenarios, where private information could also be revealed through the analysis of gradients transmitted from the devices [35, 36, 37].

### v2

@@ V Simulations
In this section, we evaluate the performance of ACFL through simulations. Although ACFL is designed and analyzed for linear regression problems, it can also be applied to various nonlinear regression problems, considering that many nonlinear problems can be transformed into linear ones. To more extensively verify the effectiveness and superiority of ACFL, we consider both linear and nonlinear regression scenarios in our simulations.
V-A Linear regression scenarios
We consider an FL setting, where each device possesses a local dataset with $m_{i}=100$ samples, $\forall i$. The devices collaboratively train a linear regression model defined by a parameter matrix $\mathbf{W}\in\mathbb{R}^{d\times o}$, where $d=10$ represents the input feature dimension of the data samples and $o=10$ denotes the output dimension of the labels. The elements in each local dataset ${\mathbf{X}}^{(i)}$} are i.i.d., drawn from a uniform distribution over the interval [1,1][-1,1]. A true parameter matrix ${{\mathbf{W}}^{\text{True}}}\in{\mathbb{R}^{d\times o}}$ is generated with elements uniformly distributed within $[0,1/30]$. The labels ${\mathbf{Y}}^{(i)}$} for each device are generated according to ${{\mathbf{Y}}^{(i)}}={{\mathbf{X}}^{(i)}}\left({{{\mathbf{W}}^{\text{True}}}+i{{\mathbf{W}}^{\text{Shift}}}}\right)$ for all $i$, where ${{\mathbf{W}}^{\text{Shift}}}\in\mathbb{R}^{d\times o}$ has elements independently drawn from the uniform distribution over $0,\sigma_{s}^{2}]$. A larger value of $\sigma_{s}^{2}$ corresponds to a greater degree of non-i.i.d.-ness in the training data across devices. The initial parameter matrix ${\mathbf{W}}_{0}$ is randomly generated with elements drawn from a uniform distribution in the interval $[0,1/30]$, independent of the true parameter matrix generation.
To evaluate the effectiveness of the adaptive policy in ACFL, we conduct a comparative analysis between ACFL and its non-adaptive counterpart (denoted as NA) with $N=100$ devices, which employs a fixed aggregation weight $\alpha_{t}=0.5$, as suggested by [31]. In Fig. 5, we depict the training loss as a function of number of iterations for both ACFL and NA under different variances of additive noise [^2] Note that under the same variances of additive noise, the privacy level is the same for both methods., where we set $\sigma_{s}^{2}=0$, $\eta_{t}=0.0001/t$ and we consider two distinct scenarios with varying straggler probabilities. As anticipated, an increase in noise variances, aimed at enhancing privacy, inversely affects the learning performance of both methods. This empirical observation aligns with our theoretical analysis, underscoring the intrinsic trade-off between privacy preservation and learning performance. Notably, the performance deterioration in ACFL remains relatively modest in comparison to NA upon increasing the noise, while there is a pronounced decline in the learning performance of NA. This discrepancy demonstrates the superior capability of ACFL to leverage gradients from both the central server and devices more effectively with the adaptive policy and validates the advantage of employing the adaptive policy in ACFL for mitigating the adverse effects of privacy-enhancing measures on learning performance.
@@ V-A Linear regression scenarios
To further validate the efficacy of the proposed method, we compare its performance with the state-of-the-art method SCFL. Note that the difference between SCFL and ACFL lies in the way of encoding local datasets and the way of aggregating gradients at the central server, which has been detailed in Section I. In Fig. 6, we plot the training loss as a function of number of iterations for both methods under various privacy levels and varying degrees of non-i.i.d.-ness in the training data, i.e., under varying values of $\epsilon$ using the metric of $\epsilon$-MI-DP and varying values of $\sigma^{2}_{s}$. The variances of the noise in both methods are calculated according to Theorem 1 from this paper and Theorem 2 in [31], ensuring required value of $\epsilon$ for both methods. Here, we fix $p=0.2,N=100$ and set the learning rate as $\eta_{t}=0.0001/t$ throughout the iterations. For an equitable comparison, we conduct full-batch gradients for both methods [^3] It is important to note that the analyses and simulations presented in this paper can be readily extended to scenarios involving stochastic gradient computations by the devices and the central server.. Furthermore, the local coded datasets in both methods are of the same size. From Fig. 6, it is evident that ACFL achieves superior learning performance under the same privacy level, both in the i.i.d. setting (with $\sigma_{s}^{2}=0$) and in the non-i.i.d. setting (with $\sigma_{s}^{2}=0.001$), where the non-i.i.d. nature of the data is controlled by the value of $\sigma_{s}^{2}$. Moreover, the learning performance of ACFL exhibits greater resilience and suffers less from increasing the noise variances to heighten the privacy level. This advantage stems from two aspects. First, in ACFL, before the training, the transformed local datasets retain information which is totally equivalent to the raw datasets for gradient computation, contributing to better learning performance at the same privacy level. Second, during the training, SCFL consistently utilizes fixed aggregation weights across iterations at the server, while ACFL uses adaptive aggregation weights to aggregate information at the server by fully considering the generation process of the global coded dataset and the dynamic nature of the trained model over iterations. In this way, ACFL guarantees the optimal performance in terms of both privacy and learning.
@@
Figure 6: Training loss as a function of number of iterations of ACFL and SCFL under varying privacy levels and varying degrees of non-i.i.d.-ness in the training data. (a) $\sigma_{s}^{2}=0$. (b) $\sigma_{s}^{2}=0.001$.
Figure 7: Training loss as a function of the number of iterations for ACFL and IS with varying numbers of devices. (a) The training loss is shown on a linear scale. (b) The training loss is shown on a logarithmic scale.
Finally, we compare the proposed method with the FL method that simply ignores the existence of stragglers and uses only the information sent by the non-stragglers during the training iterations (denoted as IS). IS is equivalent to fixing the parameter as $\alpha_{t}=0$ in the proposed method. In Fig. 7, we show the training loss as a function of the number of iterations for both the proposed method and IS, with varying numbers of devices, where we fix $\sigma_{s}^{2}=0$, $p=0.8$, $\sigma_{1}=\sigma_{2}=0.2$, and set the learning rate as $\eta_{t}=0.01/t$. We can observe in Fig. 7 that, during the early stages of training, IS experiences very unstable performance due to missing information from the stragglers. In contrast, the proposed method converges stably throughout the training process, benefiting from the fact that the missing information from the stragglers can be compensated by the information gained from the global coded dataset. It is worth noting that the performance gain of ACFL comes at the cost of some privacy leakage compared to IS. In practice, a proper trade-off between learning performance and privacy preservation should be considered when applying ACFL. It is worth emphasizing that the value of our work lies in providing more flexibility to adjust the trade-off between privacy level and straggler mitigation, making the designed FL system more adaptable to different scenarios.
V-B Nonlinear regression scenarios
In practice, many nonlinear problems can be transformed into linear regression problems, indicating that the proposed method has broader applications beyond linear cases. To verify the value of our proposed method in nonlinear problems, we consider an FL setting with nonlinear training loss, where each device holds a local dataset with $m_{i}=100$ samples, $\forall i$. The local dataset at device $i$ is represented by ${{\mathbf{x}}^{(i)}}\in\mathbb{R}^{m_{i}\times 1}$, where $x^{i}_{k}\in\mathbb{R}$ denotes the $k$-th data sample in ${\mathbf{x}}^{(i)}$}, $k=1,...,m_{i}$. The elements in ${\mathbf{x}}^{(i)}$} are i.i.d., drawn from the uniform distribution over the interval [1,1][-1,1]. The label ${{y}}^{i}_{k}$} corresponding to $x^{i}_{k}$ is generated according to $y_{k}^{i}=\exp\left({\xi x_{k}^{i}}\right)-1$, $\forall i,k$, where $\exp\left({\cdot}\right)$ is the exponential function, and $\xi=0.5$ is a constant. To fit the nonlinear relationship between the data samples and the labels, we use a high-order approximation by expanding each training data sample as $\left[{x_{k}^{i},{{\left({x_{k}^{i}}\right)}^{2}},...,{{\left({x_{k}^{i}}\right)}^{10}}}\right]$ and applying the linear regression model expressed as ${{\mathbf{y}}^{\left(i\right)}}={{\mathbf{X}}^{\left(i\right)}}{\mathbf{w}}$, where ${{\mathbf{y}}^{(i)}}\in\mathbb{R}^{m_{i}\times 1}$ contains all the labels of the local dataset for device $i$. Here, ${{\mathbf{X}}^{(i)}}\in\mathbb{R}^{m_{i}\times 10}$ represents the expanded local dataset of device $i$, where the $k$-th row of ${\mathbf{X}}^{(i)}$} is given as $\left[{x_{k}^{i},{{\left({x_{k}^{i}}\right)}^{2}},...,{{\left({x_{k}^{i}}\right)}^{10}}}\right]$, $\forall k$. The vector $\mathbf{w}\in\mathbb{R}^{10\times 1}$ is the model parameter to be optimized, which indicates that the devices collaboratively train a linear regression model defined by the vector $\mathbf{w}\in\mathbb{R}^{d\times o}$, where $d=10$ represents the input feature dimension and $o=1$ denotes the output dimension. Before the training, the initial parameter vector ${\mathbf{w}}_{0}$ is generated, where the elements are drawn from a uniform distribution in the interval $[0,1]$. To test the performance of the trained model, a test set is generated consisting of 10 data samples drawn from the same distribution as the training data. The corresponding labels are generated in the same manner as those in the training set.
Figure 8: Training loss and test loss as functions of the number of iterations for ACFL and SCFL under varying privacy levels, where each device holds a local dataset with $m_{i}=100$ samples, $\forall i$. (a) The training loss. (b) The test loss.
To verify the superiority of the proposed method, we compare its performance with SCFL by plotting the training loss and test loss as functions of the number of iterations under various privacy levels in Fig. 8. Here, we fix $p=0.2$, set $N=100$, and use a learning rate of $\eta_{t}=0.0001/t$ during training. From Fig. 8a, we can clearly observe that ACFL achieves significantly better learning performance at the same privacy level, attaining a lower training loss after the same number of iterations and converging faster. Similarly, Fig. 8b shows that ACFL also outperforms SCFL in terms of test performance. These results demonstrate that the proposed method holds significant value in FL under nonlinear regression scenarios.
To further demonstrate the superiority of the proposed method when each device possesses more data samples, we compare the performance of ACFL with SCFL by plotting the training and test losses as functions of the number of iterations under various privacy levels in Fig. 9. In this setup, each device holds a local dataset with $m_{i}=1000$ samples; we set $p=0.2$, $N=100$, and use a learning rate of $\eta_{t}=0.00001/t$. From Fig. 9, it can be observed that ACFL achieves significantly better learning performance under the same privacy level, attaining lower training and test losses after the same number of iterations. These results indicate that ACFL outperforms the baseline across various settings, regardless of whether each device has a smaller or larger amount of training data in FL under nonlinear regression scenarios.
Figure 9: Training loss and test loss as functions of the number of iterations for ACFL and SCFL under varying privacy levels, where each device holds a local dataset with $m_{i}=1000$ samples, $\forall i$. (a) The training loss. (b) The test loss.
In this simulation, we specifically consider a simple nonlinear regression model based on the power law relationship. This choice is motivated by the fact that power law models have widespread applications, as numerous physical and human-made phenomena approximately follow power law relationships. Examples include modeling in the human motor system [42] and the modeling of banking networks [43]. It is worth emphasizing that the proposed method can also be applied to other nonlinear regression models using various techniques that transform nonlinear regression problems into linear ones. However, due to page limitations, extensive examples and simulations are not provided in this paper. Investigating scalability and practical performance in more complex scenarios will be considered as a promising direction for future research.
@@ VI Conclusions
In this article, we addressed the problem of FL in the presence of stragglers. For this issue, the CFL paradigm was explored in the literature. However, the aggregation of gradients at the central server in existing CFL methods, which employ fixed weights across iterations, leads to diminished learning performance. To overcome this limitation, we proposed a new method: ACFL. In ACFL, each device uploads a local coded dataset to the central server before training, enabling the central server to generate a global coded dataset under privacy-preservation requirements. During each iteration of the training, the central server aggregates the received gradients as well as the gradient computed from the global coded dataset by using an adaptive policy to determine the aggregation weights. This approach optimizes the performance in terms of privacy and learning under theoretical guarantees. Finally, we provided simulation results to demonstrate the superiority of ACFL compared to a number of state-of-the-art baseline methods.
It is worth noting that in this work, we focused on the privacy preservation of raw datasets of the devices. In the future, we plan to extend the proposed method to more complex scenarios, where private information could also be revealed through the analysis of gradients transmitted from the devices and the devices do not share raw local gradients with the server [44, 45, 46]. Additionally, we plan to propose an enhanced version of the ACFL scheme by allowing devices to perform varying amounts of work in each iteration based on their available resources or using partial computations of the stragglers, rather than directly dropping the stragglers [47]. We would also like to point out that, in this work, we assumed that the server is trustworthy and operates honestly, as has been similarly assumed in most of the existing FL literature. In the future, we plan to extend this work to more complex environments where the server may be malicious based on verifiable FL techniques [48]. In future work, we could also explore how multi-modal learning can preserve privacy under the CFL framework by leveraging modality-specific noise injection [6].
Appendix A Proof of Theorem 1
$\mathop{\sup}\limits_{j,k,\Gamma\left({{\mathbf{X}},{\mathbf{Y}}}\right)}I\left({\left.{X_{j,k};{\mathbf{H}}_{X},{\mathbf{H}}_{Y}}\right|{{\mathbf{X}}^{-j,k}},{\mathbf{Y}}}\right)$
$=$ $\mathop{\sup}\limits_{\Gamma\left({{\mathbf{X}},{\mathbf{Y}}}\right)}I\left({\left.{X_{1,1};{\mathbf{H}}_{X},{\mathbf{H}}_{Y}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right),$ (35)
it is sufficient to show
$\mathop{\sup}\limits_{\Gamma\left({{\mathbf{X}},{\mathbf{Y}}}\right)}I\left({\left.{X_{1,1};{\mathbf{H}}_{X},{\mathbf{H}}_{Y}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)\leq\epsilon,$ (36)
$\mathop{\sup}\limits_{\Gamma\left({{\mathbf{X}},{\mathbf{Y}}}\right)}I\left({\left.{X_{1,1};{\mathbf{H}}_{X},{\mathbf{H}}_{Y}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)$
$=$ $\mathop{\sup}\limits_{\Gamma\left({{\mathbf{X}},{\mathbf{Y}}}\right)}I\left({\left.{X_{1,1};{{\mathbf{X}}^{T}}{\mathbf{X}}+{\mathbf{N}}_{1},{{\mathbf{X}}^{T}}{\mathbf{Y}}+{\mathbf{N}}_{2}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)$
$=$ $\mathop{\sup}\limits_{\Gamma\left({{\mathbf{X}},{\mathbf{Y}}}\right)}h\left({\left.{{{\mathbf{X}}^{T}}{\mathbf{X}}+{\mathbf{N}}_{1},{{\mathbf{X}}^{T}}{\mathbf{Y}}+{\mathbf{N}}_{2}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)$
$=$ $\mathop{\sup}\limits_{\Gamma\left({{\mathbf{X}},{\mathbf{Y}}}\right)}h\left({\left.{{{\mathbf{X}}^{T}}{\mathbf{X}}+{\mathbf{N}}_{1},{{\mathbf{X}}^{T}}{\mathbf{Y}}+{\mathbf{N}}_{2}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)$
$-h\left({{\mathbf{N}}_{1},{\mathbf{N}}_{2}}\right),$ (37)
where $h\left({\cdot}\right)$ denotes the entropy, the first equality is derived by substituting (2) and (3) into (A) and the second equality holds according to the definition of mutual information. In (A), for the first term, we have (A) based on the independence between the additive noise and the raw data.
$\mathop{\sup}\limits_{\Gamma\left({{\mathbf{X}},{\mathbf{Y}}}\right)}h\left({\left.{{{\mathbf{X}}^{T}}{\mathbf{X}}+{\mathbf{N}}_{1},{{\mathbf{X}}^{T}}{\mathbf{Y}}+{\mathbf{N}}_{2}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)$
$=$ $\mathop{\sup}\limits_{\Gamma\left({{\mathbf{X}},{\mathbf{Y}}}\right)}h\left({\left.{\left\{{{X_{1,j}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{j,1}},j\neq 1}\right\},\left\{{{X_{1,k}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{1,k}},k\neq 1}\right\},X_{1,1}^{2}+{{\left({{N_{1}}}\right)}_{1,1}},\left\{{{X_{1,1}}{Y_{1,k}}+{{\left({{N_{2}}}\right)}_{1,k}},\forall k}\right\}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)$
$\leq$ $\mathop{\sup}\limits_{\Gamma\left({\mathbf{X}}\right)}\sum\limits_{j\neq 1}{h\left({\left.{{X_{1,j}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{j,1}}}\right|{{\mathbf{X}}^{-1,1}}}\right)}+\mathop{\sup}\limits_{\Gamma\left({\mathbf{X}}\right)}\sum\limits_{k\neq 1}{h\left({\left.{{X_{1,k}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{1,k}}}\right|{{\mathbf{X}}^{-1,1}}}\right)}+\mathop{\sup}\limits_{\Gamma\left({{X_{1,1}}}\right)}h\left({X_{1,1}^{2}+{{\left({{N_{1}}}\right)}_{1,1}}}\right)$
$+\mathop{\sup}\limits_{\Gamma\left({\mathbf{Y}}\right)}\sum\limits_{k}{h\left({\left.{{X_{1,1}}{Y_{1,k}}+{{\left({{N_{2}}}\right)}_{1,k}}}\right|{\mathbf{Y}}}\right)}+h\left({\left\{{{{\left({{{N}}_{1}}\right)}_{j,k}},j\neq 1,k\neq 1}\right\}}\right)+h\left({\left\{{{{\left({{{N}}_{2}}\right)}_{j,k}},j\neq 1,\forall k}\right\}}\right)$
$\leq$ $\mathop{\sup}\limits_{\Gamma\left({\mathbf{X}}\right)}\sum\limits_{j\neq 1}{h\left({{X_{1,j}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{j,1}}}\right)}+\mathop{\sup}\limits_{\Gamma\left({\mathbf{X}}\right)}\sum\limits_{k\neq 1}{h\left({{X_{1,k}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{1,k}}}\right)}+\mathop{\sup}\limits_{\Gamma\left({{X_{1,1}}}\right)}h\left({X_{1,1}^{2}+{{\left({{N_{1}}}\right)}_{1,1}}}\right)$
$+\mathop{\sup}\limits_{\Gamma\left({\mathbf{Y}}\right)}\sum\limits_{k}{h\left({{X_{1,1}}{Y_{1,k}}+{{\left({{N_{2}}}\right)}_{1,k}}}\right)}+h\left({\left\{{{{\left({{{N}}_{1}}\right)}_{j,k}},j\neq 1,k\neq 1}\right\}}\right)+h\left({\left\{{{{\left({{{N}}_{2}}\right)}_{j,k}},j\neq 1,\forall k}\right\}}\right).$ (38)
Substituting (A) into (A) yields
$\mathop{\sup}\limits_{\Gamma\left({{\mathbf{X}},{\mathbf{Y}}}\right)}I\left({\left.{X_{1,1};{\mathbf{H}}_{X},{\mathbf{H}}_{Y}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)$
$\leq$ $\mathop{\sup}\limits_{\Gamma\left({\mathbf{X}}\right)}\sum\limits_{j\neq 1}{h\left({{X_{1,j}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{j,1}}}\right)}$
$+\mathop{\sup}\limits_{\Gamma\left({\mathbf{X}}\right)}\sum\limits_{k\neq 1}{h\left({{X_{1,k}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{1,k}}}\right)}$
$+\mathop{\sup}\limits_{\Gamma\left({{X_{1,1}}}\right)}h\left({X_{1,1}^{2}+{{\left({{N_{1}}}\right)}_{1,1}}}\right)$
$+\mathop{\sup}\limits_{\Gamma\left({\mathbf{Y}}\right)}\sum\limits_{k}{h\left({{X_{1,1}}{Y_{1,k}}+{{\left({{N_{2}}}\right)}_{1,k}}}\right)}$
$-h\left({\left\{{{{\left({{{N}}_{1}}\right)}_{j,k}},j=1,k\neq 1}\right\}}\right)$
$-h\left({\left\{{{{\left({{{N}}_{1}}\right)}_{j,k}},j\neq 1,k=1}\right\}}\right)$
based on Assumption 1 and the maximum entropy bound [49]. From (A), we can easily derive Theorem 1.
Appendix B Proof of Lemma 2
We can derive that
based on (2), (3), (4), (8), (9) and the fact that the additive noise is independent from the straggler behaviour of the devices. For the first term in (B), we have
$\times\mathbb{E}\left[{\left({{\alpha_{t}}+\frac{{1-{\alpha_{t}}}}{{1-p}}I_{t}^{\left({{i_{1}}}\right)}}\right)\left({{\alpha_{t}}+\frac{{1-{\alpha_{t}}}}{{1-p}}I_{t}^{\left({{i_{2}}}\right)}}\right)}\right]$
and the second inequality holds due to Assumption 2. For the second term in (B), we have
$=\alpha_{t}^{2}Nd\sigma_{1}^{2}\left\|{{{\mathbf{W}}_{t}}}\right\|_{F}^{2}$
$\leq\alpha_{t}^{2}Nd\sigma_{1}^{2}{C^{2}},$ (43)
based on the independence of the additive noise on different local datasets and Assumption 3. For the third term in (B), we have
Substituting (B), (B) and (44) into (B) yields Lemma 2.